OpenStudy (anonymous):

Which of the following represents the variable in this equation? 2d + 7 = 19

OpenStudy (dayakar):

d

OpenStudy (anonymous):

ya the variable is D but if you need to solve it the answer is d=6

OpenStudy (mudslinger):

13

OpenStudy (texaschic101):

the variable is D. It is an unknown number that is replaced by the variable D. but if you have to solve for the variable : 2d + 7 = 19 -- subtract 7 from both sides 2d = 19 - 7 2d = 12 -- divide both sides by 2 d = 6

OpenStudy (mudslinger):

i messed up i did it in my head it is 6